OpenStudy (anonymous):

What is the square root of -4

OpenStudy (anonymous):

2i

OpenStudy (anonymous):

okay how sure are u Raphael??

OpenStudy (theviper):

Oh yes nice problem \(\sqrt{-4}=\sqrt{-1}\times\sqrt{4}=\sqrt{-1}\times2\) Therefore we take \(\sqrt{-1}\) as iota \(i\) Therefore \(\sqrt{-4}=2i\)

OpenStudy (anonymous):

square root of -1 is "i" The definition of the square root of -1 is i, and i^2 ( "i squared") is negative 1.

OpenStudy (anonymous):

You are a genius @TheViper

OpenStudy (anonymous):

I'll second that...